OpenStudy (anonymous):

conjugate axis on the x-axis, one focus at (0, √13), equation of one directrix y= 9√13 / 13. HELP PLEASE

OpenStudy (anonymous):

what you need equation of parabola?

OpenStudy (anonymous):

hyperbola..

OpenStudy (anonymous):

since x- axis is your conjugate you get the two conics one on top and one in the bottom

OpenStudy (anonymous):

given eccentricity,e=9/sqrt13 b^2=a^2(e^2-1)

OpenStudy (anonymous):

no background about conics.. im sorry

OpenStudy (anonymous):

standard form is (x/a)^2-(y/b)^2=1 therefore your equation will be of the form (bx)^2-(ay)^2=(ab)^2

OpenStudy (anonymous):

yes yes..

OpenStudy (anonymous):

foci: (h, k + c), (h, k - c) c=sqrt(a^2+b^2) (h,k) will be your vertex equation is (y-k)^2/a^2-((x-h)/b)^2=1

OpenStudy (anonymous):

there is a change in slope (negative ) because x axis is the conjugate axis

OpenStudy (anonymous):

it would be imparted with parabola. i think?